OpenStudy (anonymous):

radical 3(7-radical3) multiply and simplify please!

OpenStudy (anonymous):



OpenStudy (anonymous):

Distribute root(3) to the parenthesis: Use the formula: \[\sqrt{a}(b - \sqrt{a}) = b \sqrt{a} - a\] Here, a = 3 and b= 7 Use this formula to solve..

OpenStudy (anonymous):

@nique504 try to solve buddy..

OpenStudy (anonymous):